OpenStudy (anonymous):

3^2x+3^(x+2) = 3_3^1

Directrix (directrix):

3^2x+3^(x+2) = 3_3^1 --> What does the dash mean here?

OpenStudy (anonymous):

3^2x+3^(x+2) = 10÷3